OpenStudy (anonymous):

Two 6-sided dice are rolled. Find the probability that exactly 1 die shows a 5. Thank You!

OpenStudy (anonymous):

What is the odd that one dice is a five when one die is rolled?

OpenStudy (anonymous):

I think its what is the odd that only one dice is a five when two are rolled

OpenStudy (anonymous):

I know but if you know the odds of one die there is a trick to find with 2

OpenStudy (anonymous):

2/12 or 1/6 probability?

OpenStudy (anonymous):

One die is 1/6. You have 2 dice so you have twice the odds. Twice the odds are 2/6 or 1/3. So the odds of one 5 in 2 dice is 1/3.

OpenStudy (anonymous):

Wouldn't twice the odds be 2/12?

OpenStudy (anonymous):

No, 1/6 and 2/12 are the same/equivalent. To double you keep the same denominator and double the numerator.

OpenStudy (phi):

You could label the dice A and B , and make a table showing all possible rolls of the two dice. then find how many of those rolls has exactly one 5.

OpenStudy (anonymous):

Thanks for the tip @Phil1111

OpenStudy (phi):

then divide by the number of possible rolls (which will be 36)

OpenStudy (anonymous):

No problem saoline. Phi great answer.

OpenStudy (welshfella):

odds of getting a 5 with one die is 1/6 but that could show on either of the die so we add the odds required probability = 1/6 + 1/6 = 1/3

OpenStudy (welshfella):

when its a matter of OR with probabilities you add the probabilities here it is a 5 on die A OR a 5 on die B = 1/6 + 1/6

OpenStudy (anonymous):

Oh okay. Thanks so much!

OpenStudy (welshfella):

yw when you see an AND - its multiply probabilities eg throw 1 die to get 5 P = 1/6 throw secodnd die to get 5 P = 1/6 P( 5 AND another 5) = 1/6 * 1/6 = 1/36

OpenStudy (anonymous):

Thank you!

OpenStudy (phi):

Find the probability that exactly 1 die shows a 5. Here is the table  there are 10 events where exactly one five shows. there are 36 equally likely possibilities \[ \frac{10}{36}= \frac{5}{18} \]